A cow is tied to one of the corners of the square field who side is 55 meters long.If the length of the rope is a tenth of the side field, what is the area of the field that is not grazed in cm2?

1
Expert's answer
2022-04-26T10:40:18-0400


The area of the field is 55 x 55= 3025 sq m.


Inside the field, the cow can only graze an area of a quarter circle with the radius of

55 / 10 = 5.5 m.


That area is (pi x radius squared) / 4, in this case approximately 23.75 sq m. because

"S = \frac{3.14 * 5.5^2} 4 = 23.75"

Subtracting 23.75 from 3025 we get 3001.25 sq m, which is the area of the field the cow can not graze.


Answer: The cow cannot graze approximately 3001.25 sq m of the square field or 30012500 sq cm.